OpenStudy (anonymous):

Differentiate the function. Then find an equation of the tangent line at the indicated point on the graph of the function.

OpenStudy (anonymous):

\[f(x)=6+\sqrt{1-x}\] at (-8,9)

OpenStudy (phi):

do you know that you can write the square root as an exponent to the 1/2 power? \[ f(x)= 6 + (1-x)^{\frac{1}{2} }\] you can use the "power rule" to take the derivative.

OpenStudy (phi):

\[ \frac{d}{dx}u^n = n u^{n-1} \frac{du}{dx}\]

OpenStudy (anonymous):

you lost me

OpenStudy (phi):

do you know the derivative of \[ x^2\] ?

OpenStudy (anonymous):

2

OpenStudy (phi):

what is the derivative of 2x ?

OpenStudy (anonymous):

2...

OpenStudy (phi):

yes the derivative of 2x is 2 the derivative of x^2 is 2x if you have time, watch this http://www.khanacademy.org/math/calculus/differential-calculus/power_rule_tutorial/v/power-rule it will be easier to explain after you watch

OpenStudy (anonymous):

I understand simple derivatives like that. I just don't know what to do with this problem

OpenStudy (phi):

what is the derivative of x^3 ?

OpenStudy (anonymous):

3x^2

OpenStudy (phi):

what is the derivative of \[ x^{\frac{1}{2} } \]?

OpenStudy (anonymous):

1/2x?

OpenStudy (phi):

same rule: multiply by the 1/2 and change the exponent by subtracting 1 1/2 -1 is the new exponent. try again.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }x ^{-\frac{ 1 }{ 2 }}\]

OpenStudy (phi):

next, what is the derivative of (x+1)^2 ?

OpenStudy (anonymous):

2x?

OpenStudy (phi):

we use the "chain rule" to take the derivative of (x+1)^2 the first step is let the problem be u^2 (where u is x+1) the derivative of u^2 is 2u u is x+1, so 2u is 2(x+1) 2nd step: after getting 2u we then multiply by du (the derivative of u) u is x+1 and du is d(x+1) = d(x) + d(1)= 1+0= 1 all that work to get du=1. we find d(x+1)^2 is 2(x+1)

OpenStudy (anonymous):

okay, then what?

OpenStudy (phi):

if you followed that, you can do your problem

OpenStudy (anonymous):

\[f ^{1}(x)=2(x+1)\]

OpenStudy (phi):

try doing \[ f(x)= 6 + (1-x)^{\frac{1}{2} } \] the derivative of the constant 6 is 0 you can try the rest...

OpenStudy (anonymous):

\[f ^{1}(x)=\frac{ 1 }{ 2 }x ^{-\frac{ 1 }{ 2 }}\]

OpenStudy (phi):

except that is the answer of the derivative of \[ f(x)= 6 + x^{\frac{1}{2} } \] you must keep the (1-x) and then multiply by the derivative of (1-x)

OpenStudy (anonymous):

I give up...

OpenStudy (anonymous):

okay, last try. \[f ^{1}(x)=\frac{ 1 }{ 2 }(1-x)^{-\frac{ 1 }{ 2 }}\]

OpenStudy (phi):

close. the last step is multiply that by the derivative of (1-x)

OpenStudy (anonymous):

what do you mean?

OpenStudy (phi):

first, what is d (1-x) ?

OpenStudy (anonymous):

-1?

OpenStudy (phi):

yes what you are doing is following this rule (see the videos on why) \[ d u^n = n u^{n-1} du \] in your problem u is (1-x) and n = 1/2 and du is d (1-x) = -1 put it all together: \[ f '(x)=-\frac{ 1 }{ 2 }(1-x)^{-\frac{ 1 }{ 2 }} \]

OpenStudy (phi):

btw, it's f prime (short for df/dx) not f^1

OpenStudy (anonymous):

okay thank you. now how do I find the equation of a tangent line?

OpenStudy (phi):

the derivative of a function represents the slope of the tangent line when you evaluate at a particular x. in your case, f'(x) evaluated at x= -8 is the slope of the tangent line at x=-8 after you find the slope, you have the slope and a point (-8,9), so you can find the equation of a line using that info.

OpenStudy (anonymous):

what's the slope?

OpenStudy (phi):

slope is what you learn about before you take calculus.

OpenStudy (anonymous):

I know what a slope is....I just don't know how to find it in these problems. I'm taking this online, am 16 and doing my best. thanks, not.

OpenStudy (phi):

in that case, you might want to watch some of Khan's videos. I posted above how to find the slope: f'(x) evaluated at x= -8 is the slope of the tangent line at x=-8 that means replace x with -8 in your derivative, do the arithmetic to simplify to a number. that number is the slope of the tangent line.

OpenStudy (anonymous):

m= -1/6

OpenStudy (anonymous):

and the tangent line is \[y=-\frac{ 1 }{ 6 }x+\frac{ 23 }{ 3 }\]

OpenStudy (phi):

yes, looks good

OpenStudy (anonymous):

Thank you!